Solve the equation 8x3 + 4x2 − 18x − 9 = 0 algebraically for all values of x.
Answer by palanisamy(496)

The given equation is
8x^3+4x^2-18x-9 = 0
8x^2(2x+1)-9(2x+1) = 0
(2x+1)(8x^2-9) = 0
2x+1 = 0 or 8x^2-9 = 0
2x=-1 or 8x^2 = 9
x = -1/2 or x^2 = 9/8
x = -1/2 or x = 3/2sqrt2 or x = -3/2sqrt2
So the solution is
x = -1/2 , 3/2sqrt2 and -3/2sqrt2